1. [7 points] Suppose that a random sample of U.S. hired farmworkers is selected from the population of these workers and tha(d) Construct the t-statistic for the slope coefficient. Is this t-statistic significant at the 10% level? Clearly show your work including the critical value which you are using.
(e) Construct the t-statistic for the intercept coefficient. Is this t-statistic significant at the 1% level? Clearly show your work including the critical value which you are using.

(f) Does least squares assumption 1 plausibly hold for this regression? Explain in detail why or why not.
(g) Are the errors in this regression plausibly homoskedastic? Explain. (Hint: It may be helpful to think of economic “stories” about what might be happening in this context for older versus younger workers in order to relate your answer to the homoskedasticity definition.)

Can you do the bottom 4, do not need help on 1,2 or 3
Answer #1

(a) (i) The slope of the regression equation is 0.081 which means that for every one year increase in the age, the average hours will increase by 0.081.

(ii) The estimated regression intercept makes no sense because, in the regression equation, age cannot be zero.

(b) (i) R-squared (R2) is a statistical measure that represents the proportion of the variance for average hours that's explained by age in a regression model.

(ii) The standard error of the estimate is a measure of the accuracy of predictions. The regression line is the line that minimizes the sum of squared deviations of prediction (also called the sum of squares error), and the standard error of the estimate is the square root of the average squared deviation.

(c) We have:

hours = 40.237 + 0.081*age

Put age = 20 years

hours = 40.237 + 0.081*20

hours = 41.857

(d) The hypothesis being tested is:

b1 = 0

b1 ≠ 0

t-statistic for the slope coefficient, t = b1/ Standard error

= 0.081/13.490 = 0.006

The p-value is 0.5023.

Since the p-value (0.5023) is greater than the significance level, we cannot reject the null hypothesis.

Therefore, we cannot conclude that this t-statistic is significant at the 10% level.

As HOMEWORKLIB answering guide, we have the option to answer only the first four sub-parts of a question in case of multiple parts. If you want me to answer the rest of the parts, please post the question in a new post. Thank You!